You have just been hired as a financial analyst for Lydex Company, a manufacturer of safety helmets. Your boss has asked you to perform a comprehensive analysis of the company’s financial statements, including comparing Lydex’s performance to its major competitors. The company’s financial statements for the last two years are as follows:

Lydex Company
Comparative Balance Sheet
This Year Last Year
Assets
Current assets:
Cash $ 1,000,000 $ 1,240,000
Marketable securities 0 300,000
Accounts receivable, net 2,860,000 1,960,000
Inventory 3,640,000 2,400,000
Prepaid expenses 270,000 210,000
Total current assets 7,770,000 6,110,000
Plant and equipment, net 9,600,000 9,090,000
Total assets $ 17,370,000 $ 15,200,000
Liabilities and Stockholders' Equity
Liabilities:
Current liabilities $ 4,050,000 $ 3,060,000
Note payable, 10% 3,700,000 3,100,000
Total liabilities 7,750,000 6,160,000
Stockholders' equity:
Common stock, $75 par value 7,500,000 7,500,000
Retained earnings 2,120,000 1,540,000
Total stockholders' equity 9,620,000 9,040,000
Total liabilities and stockholders' equity $ 17,370,000 $ 15,200,000
Lydex Company
Comparative Income Statement and Reconciliation
This Year Last Year
Sales (all on account) $ 15,900,000 $ 13,980,000
Cost of goods sold 12,720,000 10,485,000
Gross margin 3,180,000 3,495,000
Selling and administrative expenses 1,410,000 1,620,000
Net operating income 1,770,000 1,875,000
Interest expense 370,000 310,000
Net income before taxes 1,400,000 1,565,000
Income taxes (30%) 420,000 469,500
Net income 980,000 1,095,500
Common dividends 400,000 547,750
Net income retained 580,000 547,750
Beginning retained earnings 1,540,000 992,250
Ending retained earnings $ 2,120,000 $ 1,540,000

To begin your assignment you gather the following financial data and ratios that are typical of companies in Lydex Company’s industry:

Current ratio 2.3
Acid-test ratio 1.1
Average collection period 32 days
Average sale period 60 days
Return on assets 9.7 %
Debt-to-equity ratio 0.7
Times interest earned ratio 5.7
Price-earnings ratio 10

3. You decide, finally, to assess the company’s liquidity and asset management. For both this year and last year, compute:

a. Working capital.

b. The current ratio. (Round your final answers to 2 decimal places.)

c. The acid-test ratio. (Round your final answers to 2 decimal places.)

d. The average collection period. (The accounts receivable at the beginning of last year totaled $1,710,000.) (Use 365 days in a year. Round your intermediate calculations and final answers to 2 decimal place.)

e. The average sale period. (The inventory at the beginning of last year totaled $2,070,000.) (Use 365 days in a year. Round your intermediate calculations and final answers to 2 decimal place.)

f. The operating cycle. (Round your intermediate calculations and final answers to 2 decimal place.)

g. The total asset turnover. (The total assets at the beginning of last year totaled $14,650,000.) (Round your final answers to 2 decimal places.)
Answer #1

Part A

Working capital = current assets – current liabilities

This year = 7770000-4050000 = 3720000

Last year =6110000-3060000 = 3050000

Part B

Current ratio = current assets /current liabilities

This year = 7770000/4050000 = 1.92

Last year =6110000/3060000 = 2.00

Part C

Quick ratio = (cash +marketable securities + accounts receivable)/ current liabilities

This year = (1000000+2860000)/4050000 = 0.95

Last year = (1240000+300000+1960000)/3060000 = 1.14

Part D

Average collection period = 365*average accounts receivable / sales

This year = 365*((2860000+1960000)/2)/15900000 = 55.32

Last year = 365*((1960000+1710000)/2)/13980000 = 47.91

Part E

Average sale period = 365* average inventory / cost of goods sold

This year= 365*((3640000+2400000)/2)/12720000 = 86.66

Last year = 365*((2400000+2070000)/2)/10485000 = 77.80

Part F

Operating cycle = Average collection period + Average sale period

This year = 55.32+86.66 = 141.98

Last year = 47.91+77.80 = 125.71

Part G

Total asset turnover = sales / average total assets

This year = 15900000 / ((17370000+15200000)/2) = 0.98

Last year = 13980000/((15200000+14650000)/2) = 0.94